OpenStudy (anonymous):

A side of a square patio has a length of 18 feet. What is the area of the patio? need help with this one

OpenStudy (paperbacon):

Just multiply 18 by 4 since there are four sides in a square.

OpenStudy (anonymous):

is the answer 72

OpenStudy (paperbacon):

yes

OpenStudy (anonymous):

i have another one can u help me

OpenStudy (paperbacon):

sure! :)

OpenStudy (anonymous):

A rectangular piece of felt is 29 cm long and 10 cm wide. What is the area of the piece of felt?

OpenStudy (anonymous):

is this answer 290 is what i got

OpenStudy (paperbacon):

Do you know how to get the area of a rectangle?

OpenStudy (anonymous):

no

OpenStudy (paperbacon):

to get the area of a rectangle, you multiply length times width. So for your question it would be 29 times 10 and you would get 290.

OpenStudy (anonymous):

thats what i got i got 290 so will that be the answer

OpenStudy (paperbacon):

Yes

OpenStudy (anonymous):

i need u to check another 1 4 me can u do that

OpenStudy (paperbacon):

Yes

OpenStudy (paperbacon):

I would split this one in half: 6*12 and 7*4=72+28=100 ft

OpenStudy (anonymous):

so the answer is 100 r u sure

OpenStudy (paperbacon):

Yes

OpenStudy (paperbacon):

Do you go to connections?

OpenStudy (anonymous):

how is that i thought that u would multiply 6 times 12 which is 72

OpenStudy (paperbacon):

Yes, you are correct. But then you have to multiply 7*4 for the top half and get 28. 72+28 =100

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have one last one ready